A linear elastic material is subjected to differing strains. At a strain of 0.2%, the stress is 36 MPa whereas at a strain of 0.75%, the stress is 128 MPa. a) (5 points) Determine the stress in this material at a strain of 0.58% (use linear interpolation) 5, 1. b) (5 points) what is the Youngs modulus (elastic modulus) of this material? Use proper units mce
Answer #1

5)a) When strain=0.2%=0.002, stress=36 MPa, ie. stress/strain=0.002/36*106=1.8*1010 N/m2.

Also when stress=0.75%=0.0075,stress=128 MPa, i.e again stress/strain=1.8*1010 N/m2. i.e.stress/strain =constant=Young's Modulus.

Therefore when strain=0.58%=0.0058,stress=1.8*1010*0.0058=104 MPa.

b)Young's Modulus of this material=stress/strain=1.8*1010 N/m2.